OpenStudy (anonymous):

find the new numerator when the denominator changes.. 4x+1 = blank ---- ---- 3x+6 3xy+6y

OpenStudy (anonymous):

blank=4xy+y

OpenStudy (anonymous):

(4x +1)/(3x +6) = blank/ (3xy +6y) (4x +1)/(3x +6) = blank/ [y(3x+6)] => blank = (4x+1)*[y(3x+6)]/(3x+6) blank= (4x+1)y= 4xy+y

OpenStudy (anonymous):

so the answer is 4xy+y?

OpenStudy (anonymous):

Yes...